# arXiv:2602.19846v2[math.AP] 02 Mar 2026

### v2

@@ Abstract.
It is known that uniqueness of mild solutions to the incompressible Navier–Stokes equations holds in the critical class $C([0,T);L^{n}(\mathbb{R}^{n}))$ for $n\geqslant 3$. In this paper, we prove that this result is sharp in the sense that uniqueness fails if $L^{n}(\mathbb{R}^{n})$ is replaced by some scaling critical spaces that are even slightly larger. We achieve this through a complete classification for every pair $(p,q)$ of whether uniqueness of mild solutions in the critical Besov class $C([0,T);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$ holds or not. Our non-uniqueness mechanism produces infinitely many global solutions emanating even from zero initial state, whose large-time asymptotics are governed by non-trivial stationary flow. To the best of our knowledge, such non-unique solutions provide the first examples of non-dissipative unforced Navier–Stokes flow with critical regularity.
@@
Navier–Sotkes equations, non-uniqueness, scaling critical Besov spaces
@@ 1. Introduction
In dimensions $n\geqslant 3$, Lions–Masmoudi [Lio-Mas] and Furioli–Rieusset–Terraneo [Fur-Rie-Ter-2000] proved that mild solutions to (1.1) are unique in the critical space $C([0,T);L^{n}(\mathbb{R}^{n}))$. The aim of this paper is to prove the sharpness of the choice of $L^{n}(\mathbb{R}^{n})$ in the sense that the uniqueness fails as one replaces $L^{n}(\mathbb{R}^{n})$ by a scaling critical space that is even slightly larger; for instance, enlarging $L^{n}(\mathbb{R}^{n})=\dot{F}_{n,2}^{0}(\mathbb{R}^{n})$ to a slightly larger Lizorkin–Triebel space $\dot{F}_{n,q}^{0}(\mathbb{R}^{n})$ with any $q>2$ leads to non-uniqueness of the mild solutions in $C([0,T);\dot{F}_{n,q}^{0}(\mathbb{R}^{n}))$. To achieve this, we show that the necessary and sufficient condition for the uniqueness of mild solutions to (1.1) in the scaling critical Besov spaces is given by $1\leqslant p<n$ with $1\leqslant q\leqslant\infty$, or $p=n$ with $1\leqslant q\leqslant 2$; see Figure 1 below. Moreover, for the non-unique case $p=n$ with $2<q\leqslant\infty$, or $n<p\leqslant\infty$ with $1\leqslant q\leqslant\infty$, we prove that the zero initial datum $u_{0}=0$ generates infinitely many small global solutions $u\in C([0,\infty);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$, which asymptotically converge to a non-trivial flow $U\in\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ solving the unforced stationary Navier–Stokes equations:
@@ 1.1. Known results related to our work
with $1\leqslant p<\infty$ and $1\leqslant q\leqslant\infty$. This line of research originates from the pioneering work of Fujita–Kato [Fuj-Kato-64-ARMA], where they proved well-posedness in $\dot{H}^{n/2-1}(\mathbb{R}^{n})$, and was extended to $L^{n}(\mathbb{R}^{n})$ by Kato [Kato-84] and Giga–Miyakawa [Gig-Miy-85]. Moreover, Kozono–Yamazaki [Koz-Yam-94], Cannone–Planchon [Can-Pla-96], and Planchon [Pla-98] proved well-posedness in $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ for $1\leqslant p<\infty$ and $1\leqslant q\leqslant\infty$. In the endpoint case $p=\infty$, Koch–Tataru [Koh-Tat-01] established well-posedness for small data in ${\rm BMO}^{-1}(\mathbb{R}^{n})$, however it was shown by [Bou-Pav-08, Yon-10, Wan-15] that (1.1) is ill-posed in $\dot{B}_{\infty,q}^{-1}(\mathbb{R}^{n})$ ($1\leqslant q\leqslant\infty$). See [IN] for more precise analysis on the well-posedness of (1.1) in critical Besov or Lizorkin–Triebel spaces. We refer to [Kozono-Shimizu-2018, Takeuchi-2025-SIMA, Yam-2000, arXiv:2509.21272v2] for the forced Navier–Stokes flow in critical spaces.
@@
Finally, we briefly review works on non-uniqueness. One major direction originates from applying the Nash convex integration to the Euler equations on the torus (see, [DeLellis-Szekelyhidi-09, Isett-18, Daneri-Szekelyhidi-2017-ARMA, Lellis-Szekelyhidi-2013-Invent, Buckmaster-Lellis-Isett-Szekelyhidi-2015-AnnMath] for instance). Using an iterative scheme based on the Euler–Reynolds system via the geometric lemma, these works constructed Euler flows violating the energy conservation law and made substantial progress toward the Onsager conjecture. Although incorporating the viscous term was long thought to be fundamentally difficult, Buckmaster–Vicol [Buc-Vic-19] introduced the concept of intermittent convex integration and proved that uniqueness of weak solutions to the Navier–Stokes equations on $\mathbb{T}^{3}$ breaks down in the supercritical class $C([0,T);H^{\beta}(\mathbb{T}^{3}))$ for some $0<\beta\ll 1/2$. Cheskidov–Luo [Che-Luo-21] showed that an endpoint Ladyzhenskaya–Prodi–Serrin class $L^{2}(0,T;L^{\infty}(\mathbb{T}^{n}))$, is optimal with respect to the time integrability exponent; they constructed non-unique solutions in $L^{p}(0,T;L^{\infty}(\mathbb{T}^{3}))$ for $p<2$. In [Che-Luo-22-AnnPDE], they also considered another endpoint in two dimensions and proved non-uniqueness in $C([0,T);L^{p}(\mathbb{T}^{2}))$ for $p<2$. Furthermore, Cheskidov–Zeng–Zhang [arXiv:2503.05692] claimed that every initial datum in $H^{1/2}(\mathbb{T}^{3})$ generates infinitely many Navier–Stokes flows whose kinetic energy is continuous and monotonically decreasing in time. Remarkably, Coiculescu–Palasek [Coi-Pal-25] showed that even in ${\rm BMO}^{-1}(\mathbb{T}^{3})$, where well-posedness for small data is known, a certain large initial datum generates two smooth solutions. Their argument builds on ideas developed by Palasek [Pal-25-IMRN] for the dyadic Navier–Stokes model. Pushing this direction further, Cheskidov–Dai–Palasek [arXiv:2511.09556] observed that non-uniqueness, in the form of instantaneous Type I blow-up, may be generated from arbitrary $C^{\infty}$ solenoidal initial data. Another strategy to prove non-uniqueness is developed by Jia–Sverák [Jia-Sve-Invent, Jia-Sve-JFA], and is to exploit the instability of certain self-similar solutions. A notable result in this direction is due to Albritton–Brué–Colombo [Alb-Bru-Col-AnnMath], who proved non-uniqueness of Leray–Hopf weak solutions to the forced Navier–Stokes equations. See Aoki–Maekawa [Aok-Mae] for the corresponding result on a two-dimensional half space. More recently, Hou–Wang–Yang [2509.25116v1] announced an analogous statement in three-dimension with the unforced case $f\equiv 0$ by a computer-assisted proof.
@@ 1.2. Main results
Then, there exist a monotone decreasing sequence $\{\varepsilon_{m}\}_{m\in\mathbb{N}}\subset(0,1)$ with $\varepsilon_{m}\to 0$ and a constant $\eta=\eta(n,p,q)\geqslant 0$ such that for any $m\in\mathbb{N}$ and initial datum $u_{0}\in\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ with
@@
    Chemin [Che-92] and Planchon [Pla-98] established well-posedness for (1.1) in critical Besov spaces for all $1\leqslant p<\infty$ and $1\leqslant q\leqslant\infty$. This does not contradict our result, since their solution space is based on
    $C([0,T);\dot{B}_{p,q}^{\frac{n}{p}-1}(\mathbb{R}^{n}))\cap\widetilde{L^{1}}(0,T;\dot{B}_{p,q}^{\frac{n}{p}+1}(\mathbb{R}^{n})).$ (1.9)
@@
    Concerning non-uniqueness for steady Navier–Stokes flows, Luo [Luo-ARMA-2019] constructed nontrivial solutions to (1.2) belonging to $L^{2}(\mathbb{T}^{n})$ with $n\geqslant 4$. See [Rieusset-25-JFA] for two-dimensional case in $H^{-1}(\mathbb{T}^{2})$. Note that both spaces are supercritical and the three-dimensional case has been open. Our result appears to be the first to establish non-uniqueness of steady solutions in scaling critical spaces. Moreover our result includes the three-dimensional case. Meanwhile, our method cannot be implemented in two dimensions. This obstruction essentially reduces to the well-known difficulty of the solvability for the two-dimensional stationary Navier–Stokes equations on $\mathbb{R}^{2}$.
@@ 1.3. Outline of the proof
Let us explain the idea of the proof of Theorem 1.2. The strategy is that we first establish a dominant part of the solution which lies in $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ with $(p,q)$ satisfying (N1) or (N2), and then solve the remaining perturbative problem by a fixed point argument in a function space where well-posedness is available. More precisely, we construct an almost frequency-localized building blocks $\{V_{j}\}_{j\geqslant 0}$ of stationary flows, each of which are essentially built on the $j$th frequency shell, and design the principal profile
@@
belongs to well-posedness class $\dot{B}_{r,1}^{n/r-3}(\mathbb{R}^{n})$ ($r<n$) and admits the following smallness hierarchy:
@@
Furthermore, to construct a non-stationary solution $u$ whose asymptotic profile is $U$, we decompose $u=U+w$. Then, we may follow the standard stability argument to construct a small dissipative perturbed flow $w\in C([0,\infty);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))\cap\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{n/p-1+2/\theta}(\mathbb{R}^{n}))$ with some $2<\theta<\infty$. Then, the auxilliary dissipative space enables us to show $\|w(t)\|_{\dot{B}_{p,q}^{n/p-1}}\to 0$ as $t\to\infty$.
We now explain how to construct the leading profile $V$ satisfying (1.13). A natural first attempt, motivated by existing works such as [Coi-Pal-25], would be to choose each $V_{j}$ as a superposition of plane waves concentrated on a single highly lacunary frequency scale $N_{j}=A^{\gamma b^{j}}$ ($A\gg 1$), for instance, $V_{j}(x)=\sum_{k\in\mathcal{K}}f_{k,j}(x)\cos(N_{j}a_{k}\cdot x)k$, where $\mathcal{K}\subset\mathbb{R}^{n}$ is a finite set, $f_{k,j}$ are smooth amplitudes, and $a_{k}$ are constant vectors with $a_{k}\cdot k=0$. However, such a single-scale construction is too rigid to capture the delicate dependence on the Besov summability index $q$, which is essential for observing the subtle non-uniqueness phenomena addressed in (N1). Instead, we distribute the $j$th building block over many frequency scales. To this end, for $\lambda\gg 1$ (which is a parameter corresponding to $\varepsilon^{-1/\alpha}$ with some $\alpha$) and a fixed large number $\mu\in\mathbb{N}$, we introduce a tetration-like sequence [^4] This is defined by $\lambda_{0}=\lambda$ and $\lambda_{j}=\lambda^{2^{\lambda_{j-1}^{4\mu}}}$; see (4.9). $\{\lambda_{j}\}_{j\geqslant 0}$ and define the $j$th level frequency index set by
@@
${\left\|V_{j}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\sim\sum_{k\in\mathcal{K}}\frac{{\left\|f_{k,j}\right\|}_{L^{p}}}{\sqrt{h_{j}}}\left(\sum_{j\in\Lambda_{j}}\frac{\lambda^{(\frac{n}{p}-1)q\ell}}{\ell^{\frac{q}{2}}}\right)^{\frac{1}{q}},$ (1.17)
@@
The cross-shell interaction term $F_{2}$ is favorable, as it remains in high-frequency region. By contrast, $F_{1}$ exhibits a more delicate structure. Indeed, within the nonlinear interaction $\operatorname{div}(V_{j}\otimes V_{j})$, resonances among modes inside the frequency shell $\Lambda_{j}$ generate an inverse-cascade transfer toward low frequencies, which exhibit the worst regularity. By selecting the directional set $\mathcal{K}$ and the amplitudes $f_{k,j}$ appropriately through the Nash geometric lemma, one may arrange for this low-frequency contribution to cancel the linear part $\Delta V_{j-1}$. Consequently, the forcing term $F$ attains improved regularity and smaller magnitude relative to $V$.

### v3

# Sharp non-uniqueness for the Navier–Stokes equations in scaling critical spaces

@@ Abstract.
It is known that uniqueness of mild solutions to the incompressible Navier–Stokes equations holds in the critical class $C([0,T);L^{n}(\mathbb{R}^{n}))$ for $n\geqslant 3$. In this paper, we prove that this result is sharp in the sense that uniqueness fails if $L^{n}(\mathbb{R}^{n})$ is replaced by some scaling critical spaces that are even slightly larger. We achieve this through a complete classification for every pair $(p,q)$ of whether uniqueness of mild solutions in the critical Besov class $C([0,T);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$ holds or not. Our non-uniqueness mechanism produces infinitely many global solutions emanating even from zero initial state, whose large-time asymptotics are governed by non-trivial stationary flows. To the best of our knowledge, such non-unique solutions provide the first examples of non-dissipative unforced Navier–Stokes flow with critical regularity.
@@
Navier–Stokes equations, non-uniqueness, scaling critical Besov spaces
@@ 1. Introduction
In dimensions $n\geqslant 3$, Lions–Masmoudi [Lio-Mas] and Furioli–Rieusset–Terraneo [Fur-Rie-Ter-2000] proved that mild solutions to (1.1) are unique in the critical space $C([0,T);L^{n}(\mathbb{R}^{n}))$. The aim of this paper is to prove the sharpness of the choice of $L^{n}(\mathbb{R}^{n})$ in the sense that the uniqueness fails if one replaces $L^{n}(\mathbb{R}^{n})$ by a scaling critical space that is even slightly larger; for instance, enlarging $L^{n}(\mathbb{R}^{n})=\dot{F}_{n,2}^{0}(\mathbb{R}^{n})$ to a slightly larger Lizorkin–Triebel space $\dot{F}_{n,q}^{0}(\mathbb{R}^{n})$ with any $q>2$ leads to non-uniqueness of the mild solutions in $C([0,T);\dot{F}_{n,q}^{0}(\mathbb{R}^{n}))$. To achieve this, we show that the necessary and sufficient condition for the uniqueness of mild solutions to (1.1) in the scaling critical Besov spaces is given by $1\leqslant p<n$ with $1\leqslant q\leqslant\infty$, or $p=n$ with $1\leqslant q\leqslant 2$; see Figure 1 below. Moreover, for the non-unique case $p=n$ with $2<q\leqslant\infty$, or $n<p\leqslant\infty$ with $1\leqslant q\leqslant\infty$, we prove that the zero initial datum $u_{0}=0$ generates infinitely many small global solutions $u\in C([0,\infty);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$, which asymptotically converge to a non-trivial flow $U\in\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ solving the unforced stationary Navier–Stokes equations:
@@ 1.1. Known results related to our work
with $1\leqslant p<\infty$ and $1\leqslant q\leqslant\infty$. This line of research originates from the pioneering work of Fujita–Kato [Fuj-Kato-64-ARMA], where they proved well-posedness in $\dot{H}^{n/2-1}(\mathbb{R}^{n})$, and was extended to $L^{n}(\mathbb{R}^{n})$ by Kato [Kato-84] and Giga–Miyakawa [Gig-Miy-85]. Moreover, Kozono–Yamazaki [Koz-Yam-94], Cannone–Planchon [Can-Pla-96], and Planchon [Pla-98] proved well-posedness in $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ for $1\leqslant p<\infty$ and $1\leqslant q\leqslant\infty$. In the endpoint case $p=\infty$, Koch–Tataru [Koh-Tat-01] established well-posedness for small data in ${\rm BMO}^{-1}(\mathbb{R}^{n})$, however, it was shown by [Bou-Pav-08, Yon-10, Wan-15] that (1.1) is ill-posed in $\dot{B}_{\infty,q}^{-1}(\mathbb{R}^{n})$ ($1\leqslant q\leqslant\infty$). See [IN] for more precise analysis on the well-posedness of (1.1) in critical Besov or Lizorkin–Triebel spaces. We refer to [Kozono-Shimizu-2018, Takeuchi-2025-SIMA, Yam-2000, arXiv:2509.21272v2] for the forced Navier–Stokes flow in critical spaces.
@@
Finally, we briefly review works on non-uniqueness. One major direction originates from applying the Nash convex integration to the Euler equations on the torus (see for instance [DeLellis-Szekelyhidi-09, Isett-18, Daneri-Szekelyhidi-2017-ARMA, Lellis-Szekelyhidi-2013-Invent, Buckmaster-Lellis-Isett-Szekelyhidi-2015-AnnMath]). Using an iterative scheme based on the Euler–Reynolds system via the geometric lemma, these works constructed Euler flows violating the energy conservation law and made substantial progress toward the Onsager conjecture. Although incorporating the viscous term was long thought to be fundamentally difficult, Buckmaster–Vicol [Buc-Vic-19] introduced the concept of intermittent convex integration and proved that uniqueness of weak solutions to the Navier–Stokes equations on $\mathbb{T}^{3}$ breaks down in the supercritical class $C([0,T);H^{\beta}(\mathbb{T}^{3}))$ for some $0<\beta\ll 1/2$. Cheskidov–Luo [Che-Luo-21] showed that an endpoint Ladyzhenskaya–Prodi–Serrin class $L^{2}(0,T;L^{\infty}(\mathbb{T}^{n}))$ is optimal with respect to the time integrability exponent; they constructed non-unique solutions in $L^{p}(0,T;L^{\infty}(\mathbb{T}^{3}))$ for $p<2$. In [Che-Luo-22-AnnPDE], they also considered another endpoint in two dimensions and proved non-uniqueness in $C([0,T);L^{p}(\mathbb{T}^{2}))$ for $p<2$. Furthermore, Cheskidov–Zeng–Zhang [arXiv:2503.05692] claimed that every initial datum in $H^{1/2}(\mathbb{T}^{3})$ generates infinitely many Navier–Stokes flows whose kinetic energy is continuous and monotonically decreasing in time. Remarkably, Coiculescu–Palasek [Coi-Pal-25] showed that even in ${\rm BMO}^{-1}(\mathbb{T}^{3})$, where well-posedness for small data is known, a certain large initial datum generates two smooth solutions. Their argument builds on ideas developed by Palasek [Pal-25-IMRN] for the dyadic Navier–Stokes model. Pushing this direction further, Cheskidov–Dai–Palasek [arXiv:2511.09556] observed that non-uniqueness, in the form of instantaneous Type I blow-up, may be generated from arbitrary $C^{\infty}$ solenoidal initial data. Another strategy to prove non-uniqueness is developed by Jia–Sverák [Jia-Sve-Invent, Jia-Sve-JFA], and is to exploit the instability of certain self-similar solutions. A notable result in this direction is due to Albritton–Brué–Colombo [Alb-Bru-Col-AnnMath], who proved non-uniqueness of Leray–Hopf weak solutions to the forced Navier–Stokes equations. See Aoki–Maekawa [Aok-Mae] for the corresponding result on a two-dimensional half space. More recently, Hou–Wang–Yang [2509.25116v1] announced an analogous statement in three dimensions with the unforced case $f\equiv 0$ by a computer-assisted proof.
@@ 1.2. Main results
Then, there exist a monotone decreasing sequence $\{\varepsilon_{m}\}_{m\in\mathbb{N}}\subset(0,\infty)$ with $\varepsilon_{m}\to 0$ as $m\to\infty$ and a constant [^1] We choose $\eta>0$ for $p<2n$ and $q<\infty$, while $\eta=0$ for the other cases. $\eta=\eta(n,p,q)\geqslant 0$ such that for every $m\in\mathbb{N}$ and any initial datum $u_{0}\in\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ with
@@
    Chemin [Che-92] and Planchon [Pla-98] established well-posedness for (1.1) in critical Besov spaces for all $1\leqslant p<\infty$ and $1\leqslant q\leqslant\infty$. This does not contradict our result, since their solution space is
    $C([0,T);\dot{B}_{p,q}^{\frac{n}{p}-1}(\mathbb{R}^{n}))\cap\widetilde{L^{1}}(0,T;\dot{B}_{p,q}^{\frac{n}{p}+1}(\mathbb{R}^{n})),$ (1.9)
@@
    Concerning non-uniqueness for steady Navier–Stokes flows, Luo [Luo-ARMA-2019] constructed nontrivial solutions to (1.2) belonging to $L^{2}(\mathbb{T}^{n})$ with $n\geqslant 4$. See [Rieusset-25-JFA] for two-dimensional case in $H^{-1}(\mathbb{T}^{2})$. Note that both spaces are supercritical and the three-dimensional case has been open. Our result appears to be the first to establish non-uniqueness of steady solutions in scaling critical spaces. Moreover, our result includes the three-dimensional case. Meanwhile, our method cannot be implemented in two dimensions. This obstruction essentially reduces to the well-known difficulty of the solvability for the two-dimensional stationary Navier–Stokes equations on $\mathbb{R}^{2}$.
@@ 1.3. Outline of the proof
Let us explain the idea of the proof of Theorem 1.2. The strategy is that we first establish a dominant part of the solution which lies in $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ with $(p,q)$ satisfying (N1) or (N2), and then solve the remaining perturbative problem by a fixed point argument in a function space where well-posedness is available. More precisely, we construct almost frequency-localized building blocks $\{V_{j}\}_{j\geqslant 0}$ of stationary flows, each of which are essentially built on the $j$th frequency shell, and design the principal profile
@@
belongs to well-posedness [^4] See [Kan-Koz-Shi-19] for well-posedness of the stationary Navier–Stokes equations in critical Besov spaces. class $\dot{B}_{r,1}^{n/r-3}(\mathbb{R}^{n})$ ($r<n$) and admits the following smallness hierarchy:
@@
Furthermore, to construct a non-stationary solution $u$ whose asymptotic profile is $U$, we decompose $u=U+w$. Then, we may follow the standard stability argument to construct a small dissipative perturbed flow $w\in C([0,\infty);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))\cap\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{n/p-1+2/\theta}(\mathbb{R}^{n}))$ with some $2<\theta<\infty$. Then, the auxiliary dissipative norm enables us to show $\|w(t)\|_{\dot{B}_{p,q}^{n/p-1}}\to 0$ as $t\to\infty$.
We now explain how to construct the leading profile $V$ satisfying (1.13). A natural first attempt, motivated by existing works such as [Coi-Pal-25], would be to choose each $V_{j}$ as a superposition of plane waves concentrated on a single highly lacunary frequency scale $N_{j}=A^{\gamma b^{j}}$ ($A\gg 1$), for instance, $V_{j}(x)=\sum_{k\in\mathcal{K}}f_{k,j}(x)\cos(N_{j}a_{k}\cdot x)k$, where $\mathcal{K}\subset\mathbb{R}^{n}$ is a finite set, $f_{k,j}$ are smooth amplitudes, and $a_{k}$ are constant vectors with $a_{k}\cdot k=0$. However, such a single-scale construction is too rigid to capture the delicate dependence on the Besov summability index $q$, which is essential for observing the subtle non-uniqueness phenomena addressed in (N1). Instead, we distribute the $j$th building block over many frequency scales. To this end, for $\lambda\gg 1$ (which is a parameter corresponding to $\varepsilon_{m}^{-1/\alpha}$ with some $\alpha$) and a fixed large number $\mu\in\mathbb{N}$, we introduce a tetration-like sequence [^6] This is defined by $\lambda_{0}=\lambda$ and $\lambda_{j}=\lambda^{2^{\lambda_{j-1}^{4\mu}}}$; see (4.9). $\{\lambda_{j}\}_{j\geqslant 0}$ and define the $j$th level frequency index set by
@@
${\left\|V_{j}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\sim\sum_{k\in\mathcal{K}}\frac{{\left\|f_{k,j}\right\|}_{L^{p}}}{\sqrt{h_{j}}}\left(\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(\frac{n}{p}-1)q\ell}}{\ell^{\frac{q}{2}}}\right)^{\frac{1}{q}},$ (1.17)
@@
The cross-shell interaction term $F_{2}$ is favorable, as it remains in the high-frequency region. By contrast, $F_{1}$ exhibits a more delicate structure. Indeed, within the nonlinear interaction $\operatorname{div}(V_{j}\otimes V_{j})$, resonances among modes inside the frequency shell $\Lambda_{j}$ generate an inverse-cascade transfer toward low frequencies, which exhibit the worst regularity. By selecting the directional set $\mathcal{K}$ and the amplitudes $f_{k,j}$ appropriately through the Nash geometric lemma, one may arrange for this low-frequency contribution to cancel the linear part $\Delta V_{j-1}$. Consequently, the forcing term $F$ attains improved regularity and smaller magnitude relative to $V$.